# CaseLink: Inductive Graph Learning for Legal Case Retrieval

### v1

@@ Abstract.
copyright: acmlicensedjournalyear: 2024doi: XXXXXXX.XXXXXXXconference: Proceedings of the 47th International ACM SIGIR Conference on Research and Development in Information Retrieval; July 14–18, 2024; Washington D.C., USAisbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Specialized information retrieval
@@ 1. Introduction
Generally, existing LCR models can be divided into two types: statistical models and neural network models. Statistical models, such as BM25 (Robertson and Walker, 1994), TF-IDF (Jones, 2004) and LMIR (Ponte and Croft, 2017), use term frequency in documents and corpus to measure the relevance between cases. For neural network models (Chalkidis and Kampas, 2019; Xiao et al., 2021; Abolghasemi et al., 2022; Wang, 2022; Chalkidis et al., 2020; Ma et al., 2022; Althammer et al., 2021; Askari et al., 2023b; Vuong et al., 2023; Rabelo et al., 2022; Yu et al., 2022; Sun et al., 2022; Askari and Verberne, 2021; Askari et al., 2023a, 2022; Yao et al., 2022; Tran et al., 2019; Zhang et al., 2023b; Zhong et al., 2020; Liu et al., 2022, 2023; Shao et al., 2020; Li et al., 2023; Tang et al., 2023, 2024; Qin et al., 2023; Ma et al., 2023b; Zhang et al., 2023a), the similarity between cases are obtained based on individually encoded representations of case texts. Typically, the encoding involves language models (BERT-PLI (Shao et al., 2020), SAILER (Li et al., 2023) and PromptCase (Tang et al., 2023)), or graph neural networks (Kipf and Welling, 2017) (CaseGNN (Tang et al., 2024)), to leverage the semantics in the case text.
@@ 2.1. Legal Case Retrieval
Legal case retrieval aims to retrieve relevant cases from a large database given a query legal case, belonging to query-by-document category. Recent LCR methods can be mainly divided into two types: statistical models (Jones, 2004; Robertson and Walker, 1994; Ponte and Croft, 2017) and neural network models (Chalkidis and Kampas, 2019; Xiao et al., 2021; Abolghasemi et al., 2022; Wang, 2022; Chalkidis et al., 2020; Ma et al., 2022; Althammer et al., 2021; Askari et al., 2023b; Vuong et al., 2023; Rabelo et al., 2022; Yu et al., 2022; Sun et al., 2022; Askari and Verberne, 2021; Askari et al., 2023a, 2022; Yao et al., 2022; Tran et al., 2019; Zhang et al., 2023b; Zhong et al., 2020; Liu et al., 2022, 2023; Shao et al., 2020; Li et al., 2023; Tang et al., 2023, 2024; Ma et al., 2023b; Qin et al., 2023; Zhou et al., 2023). For statistical models, most methods rely on using term frequency to measure the similarity between cases, such as TF-IDF (Jones, 2004), BM25 (Robertson and Walker, 1994) and LMIR (Ponte and Croft, 2017). These methods are convenient for conducting calculations since there is only term frequency counting without optimisation and inference of large neural network models. One of the drawbacks of these statistical models is that only using term frequency cannot effectively represent the relevance between cases given the complexity of the individual cases. For the neural network models, most of the existing work relies on encoding the case text into a high-dimensional vector with neural networks of various structures. For example, BERT (Devlin et al., 2019) is a typical encoder that generates a representation for the case text. Given that there is an input length limit (e.g., 512 tokens) for BERT while a case can generally contain thousands of words, most BERT-based LCR methods aim to bypass this restriction while maintaining the quality of the case representation. BERT-PLI (Shao et al., 2020) proposes a paragraph-level interaction module by dividing the case into multiple paragraphs and measuring the similarity between cases by aggregating the paragraph-level similarity. SAILER (Li et al., 2023) chooses to truncate the overlong case text. One recent model, Gear (Qin et al., 2023), makes use of generative retrieval with legal judge prediction, which is similar to the case reasoning generation pre-training in SAILER. IOT-Match (Yu et al., 2022) makes use of sentence embeddings and calculates the similarity with an optimal transport distance. PromptCase (Tang et al., 2023) relies on generating a summary of the legal fact and identifying the legal issue sections together with a prompt reformulation to obtain an effective input representation. CaseGNN (Tang et al., 2024) further transforms the unstructured case text into a structured case graph for each case and encodes the case with a dedicated graph neural work model. In general, most of these existing LCR models put their effort into using the case text to encode the individual cases. In contrast, the proposed CaseLink targets at the LCR problem from the perspective of exploiting the case reference information with latent connectivity of cases.
@@ 2.2. Graph Neural Networks
Graph neural networks (GNNs) are an effective tool for representation learning on graph data (Kipf and Welling, 2017; Velickovic et al., 2018; Hamilton et al., 2017). Graph convolutional network (GCN) develops a convolution operation on graph data in a transductive setting (Kipf and Welling, 2017). Graph attention network (GAT) extends the GCN layer with an attention module (Velickovic et al., 2018). GraphSAGE further designs an inductive model with neighbour sampling (Hamilton et al., 2017). Graph has been used for many applications (Zhang et al., 2021; Gao et al., 2024; Qu et al., 2021; Sun et al., 2021; Hung et al., 2017; Qiu et al., 2019, 2020a, 2020b, 2022).
@@ 4.1.3. Case-Case Edge
According to the case relationship within the case pool, it is desired to link the cases that have intrinsic connectivity to be neighbouring nodes for a more effective message passing of case nodes. Thus, the pairwise BM25 (Robertson and Walker, 1994) score between every cases is calculated. Due to the large number of cases, only top $k$ highest BM25 similarity scores cases are selected and linked as blue edges in Figure 2. Although BM25 is a non-symmetric measurement, which essentially makes the case-case edge directed, for the simplicity of CaseLink framework, the case-case edge is converted into symmetric. The final adjacency matrix of case-case edge $A_{d}\in\mathbb{R}^{n\times n}$ ($n$ cases in case pool in total) is denoted as:
(3) $A_{d_{ij}}=\left\{\begin{aligned} 1\quad&\text{for}&\text{TopK}(\text{BM25}(t_{d_{i}},t_{d_{j}}|d_{i},d_{j}\in\mathcal{D})),\\ 0\quad&\text{for}&\text{Others},\end{aligned}\right.$
@@ 4.1.4. Charge-Charge Edge
Within a legal system, there are natural relationships between different charges by considering that multiple charges may simultaneously appear in the same case. When two charges are linked by an edge, they may have a higher probabilities to both appear in two similar cases than other separated charges. The symmetric adjacency matrix of charge-charge edge $A_{c}\in\mathbb{R}^{m\times m}$ ($m$ is number of charges) defined as:
(4) $A_{c_{ij}}=\left\{\begin{aligned} 1\quad&\text{for}&\text{Sim}(\textbf{x}_{c_{i}},\textbf{x}_{c_{j}}|c_{i},c_{j}\in\mathcal{V})>\delta,\\ 0\quad&\text{for}&\text{Others},\end{aligned}\right.$
@@ 4.1.5. Case-Charge Edge
Clarifying the legal charge of a case is essential for legal practitioners to understand the case itself and identify relevant cases. The case-charge edge exists when a charge appears in a given case, which means that there is a natural one-way inclusion relationship between case and charge. The adjacency matrix of case-charge edge $A_{b}\in\mathbb{R}^{m\times n}$ is denoted as:
(5) $A_{b_{ij}}=\left\{\begin{aligned} 1\quad&\text{for}&t_{c_{i}}\ \text{appears in}\ t_{d_{j}},\\ 0\quad&\text{for}&\text{Others},\end{aligned}\right.$
@@ 4.1.6. Overall Adjacency Matrix
The edges of GCG including case-case, charge-charge and case-charge edges are undirected and unweighted. The overall adjacency matrix $A\in\mathbb{R}^{(n+m)\times(n+m)}$ is:
(6) $A=\begin{bmatrix}A_{d}&A_{b}^{T}\\ A_{b}&A_{c}\end{bmatrix},$
where $A_{b}^{T}$ is the transpose matrix of adjacency matrix $A_{b}$. The overall adjacency matrix, $A$, is symmetric.
@@ 4.3.2. Degree Regularisation of Candidate
(10) $\hat{A}_{ij}=\text{cos}(\mathbf{h}_{i},\mathbf{h}_{j}),$
where $\mathbf{h}_{i}$ and $\mathbf{h}_{j}$ are the updated features of case node $i$ and $j$ in the case pool $\mathcal{D}$. The matrix $\hat{A}\in\mathbb{R}^{n\times n}$ indicates a fully connected situation. And the degree regularisation is conducted on this pseudo adjacency matrix $\hat{A}$ only for candidate cases:
(11) $\ell_{\text{DegReg}}=\sum\limits^{o}_{i=1}\sum\limits^{n}_{j=1}(\hat{A}_{ij}),$
@@ 4.4. Inference
The overall comparison between typical case text encoding-based LCR methods and the CaseLink is demonstrated in Figure 3.

### v3

@@ Abstract.
journalyear: 2024copyright: acmlicensedconference: Proceedings of the 47th International ACM SIGIR Conference on Research and Development in Information Retrieval; July 14–18, 2024; Washington, DC, USAbooktitle: Proceedings of the 47th International ACM SIGIR Conference on Research and Development in Information Retrieval (SIGIR ’24), July 14–18, 2024, Washington, DC, USAdoi: 10.1145/3626772.3657693isbn: 979-8-4007-0431-4/24/07ccs: Information systems Specialized information retrieval
@@ 1. Introduction
Generally, existing LCR models can be divided into two types: statistical models and neural network models. Statistical models, such as BM25 (Robertson and Walker, 1994), TF-IDF (Jones, 2004) and LMIR (Ponte and Croft, 2017), use term frequency in documents and corpus to measure the relevance between cases. For neural network models (Chalkidis and Kampas, 2019; Xiao et al., 2021; Abolghasemi et al., 2022; Wang, 2022; Chalkidis et al., 2020; Ma et al., 2022; Althammer et al., 2021; Askari et al., 2023b; Vuong et al., 2023; Rabelo et al., 2022; Yu et al., 2022; Sun et al., 2022; Askari and Verberne, 2021; Askari et al., 2023a, 2022; Yao et al., 2022; Tran et al., 2019; Zhang et al., 2023b; Zhong et al., 2020; Liu et al., 2022, 2023a; Shao et al., 2020; Li et al., 2023; Tang et al., 2023, 2024; Qin et al., 2023; Ma et al., 2023b; Zhang et al., 2023a), the similarity between cases are obtained based on individually encoded representations of case texts. Typically, the encoding involves language models (BERT-PLI (Shao et al., 2020), SAILER (Li et al., 2023) and PromptCase (Tang et al., 2023)), or graph neural networks (Kipf and Welling, 2017) (CaseGNN (Tang et al., 2024)), to leverage the semantics in the case text.
@@ 2.1. Legal Case Retrieval
Legal case retrieval aims to retrieve relevant cases from a large database given a query legal case, belonging to query-by-document category. Recent LCR methods can be mainly divided into two types: statistical models (Jones, 2004; Robertson and Walker, 1994; Ponte and Croft, 2017) and neural network models (Chalkidis and Kampas, 2019; Xiao et al., 2021; Abolghasemi et al., 2022; Wang, 2022; Chalkidis et al., 2020; Ma et al., 2022; Althammer et al., 2021; Askari et al., 2023b; Vuong et al., 2023; Rabelo et al., 2022; Yu et al., 2022; Sun et al., 2022; Askari and Verberne, 2021; Askari et al., 2023a, 2022; Yao et al., 2022; Tran et al., 2019; Zhang et al., 2023b; Zhong et al., 2020; Liu et al., 2022, 2023a; Shao et al., 2020; Li et al., 2023; Tang et al., 2023, 2024; Ma et al., 2023b; Qin et al., 2023; Zhou et al., 2023). For statistical models, most methods rely on using term frequency to measure the similarity between cases, such as TF-IDF (Jones, 2004), BM25 (Robertson and Walker, 1994) and LMIR (Ponte and Croft, 2017). These methods are convenient for conducting calculations since there is only term frequency counting without optimisation and inference of large neural network models. One of the drawbacks of these statistical models is that only using term frequency cannot effectively represent the relevance between cases given the complexity of the individual cases. For the neural network models, most of the existing work relies on encoding the case text into a high-dimensional vector with neural networks of various structures. For example, BERT (Devlin et al., 2019) is a typical encoder that generates a representation for the case text. Given that there is an input length limit (e.g., 512 tokens) for BERT while a case can generally contain thousands of words, most BERT-based LCR methods aim to bypass this restriction while maintaining the quality of the case representation. BERT-PLI (Shao et al., 2020) proposes a paragraph-level interaction module by dividing the case into multiple paragraphs and measuring the similarity between cases by aggregating the paragraph-level similarity. SAILER (Li et al., 2023) chooses to truncate the overlong case text. One recent model, Gear (Qin et al., 2023), makes use of generative retrieval with legal judge prediction, which is similar to the case reasoning generation pre-training in SAILER. IOT-Match (Yu et al., 2022) makes use of sentence embeddings and calculates the similarity with an optimal transport distance. PromptCase (Tang et al., 2023) relies on generating a summary of the legal fact and identifying the legal issue sections together with a prompt reformulation to obtain an effective input representation. CaseGNN (Tang et al., 2024) further transforms the unstructured case text into a structured case graph for each case and encodes the case with a dedicated graph neural work model. In general, most of these existing LCR models put their effort into using the case text to encode the individual cases. In contrast, the proposed CaseLink targets at the LCR problem from the perspective of exploiting the case reference information with latent connectivity of cases.
@@ 2.2. Graph Neural Networks
Graph neural networks (GNNs) are an effective tool for representation learning on graph data (Kipf and Welling, 2017; Velickovic et al., 2018; Hamilton et al., 2017). Graph convolutional network (GCN) develops a convolution operation on graph data in a transductive setting (Kipf and Welling, 2017). Graph attention network (GAT) extends the GCN layer with an attention module (Velickovic et al., 2018). GraphSAGE further designs an inductive model with neighbour sampling (Hamilton et al., 2017). Graph has been used for many applications (Zhang et al., 2021; Gao et al., 2024; Qu et al., 2021; Sun et al., 2021; Hung et al., 2017; Qiu et al., 2019, 2020a, 2020b, 2022; Liu et al., 2023b, c).
@@ 4.1.3. Case-Case Edge
According to the case relationship within the case pool, it is desired to link the cases that have intrinsic connectivity to be neighbouring nodes for a more effective message passing of case nodes. Thus, the pairwise BM25 (Robertson and Walker, 1994) score between every cases is calculated. Due to the large number of cases, only top $k$ highest BM25 similarity scores cases are selected and linked as blue edges in Figure 2. Although BM25 is a non-symmetric measurement, which essentially makes the case-case edge directed, for the simplicity of CaseLink framework, the case-case edge is converted into symmetric. The final adjacency matrix of case-case edge $\textbf{A}_{d}\in\mathbb{R}^{n\times n}$ ($n$ cases in case pool in total) is denoted as:
(3) $\textbf{A}_{d_{ij}}=\left\{\begin{aligned} 1\quad&\text{for}&\text{TopK}(\text{BM25}(t_{d_{i}},t_{d_{j}}|d_{i},d_{j}\in\mathcal{D})),\\ 0\quad&\text{for}&\text{Others},\end{aligned}\right.$
@@ 4.1.4. Charge-Charge Edge
Within a legal system, there are natural relationships between different charges by considering that multiple charges may simultaneously appear in the same case. When two charges are linked by an edge, they may have a higher probabilities to both appear in two similar cases than other separated charges. The symmetric adjacency matrix of charge-charge edge $\textbf{A}_{c}\in\mathbb{R}^{m\times m}$ ($m$ is number of charges) defined as:
(4) $\textbf{A}_{c_{ij}}=\left\{\begin{aligned} 1\quad&\text{for}&\text{Sim}(\textbf{x}_{c_{i}},\textbf{x}_{c_{j}}|c_{i},c_{j}\in\mathcal{V})>\delta,\\ 0\quad&\text{for}&\text{Others},\end{aligned}\right.$
@@ 4.1.5. Case-Charge Edge
Clarifying the legal charge of a case is essential for legal practitioners to understand the case itself and identify relevant cases. The case-charge edge exists when a charge appears in a given case, which means that there is a natural one-way inclusion relationship between case and charge. The adjacency matrix of case-charge edge $\textbf{A}_{b}\in\mathbb{R}^{m\times n}$ is denoted as:
(5) $\textbf{A}_{b_{ij}}=\left\{\begin{aligned} 1\quad&\text{for}&t_{c_{i}}\ \text{appears in}\ t_{d_{j}},\\ 0\quad&\text{for}&\text{Others},\end{aligned}\right.$
@@ 4.1.6. Overall Adjacency Matrix
The edges of GCG including case-case, charge-charge and case-charge edges are undirected and unweighted. The overall adjacency matrix $\textbf{A}\in\mathbb{R}^{(n+m)\times(n+m)}$ is:
(6) $\textbf{A}=\begin{bmatrix}\textbf{A}_{d}&\textbf{A}_{b}^{T}\\ \textbf{A}_{b}&\textbf{A}_{c}\end{bmatrix},$
where $\textbf{A}_{b}^{T}$ is the transpose matrix of adjacency matrix $\textbf{A}_{b}$. The overall adjacency matrix, A, is symmetric.
@@ 4.3.2. Degree Regularisation of Candidate
(10) $\hat{\textbf{A}}_{ij}=\text{cos}(\mathbf{h}_{i},\mathbf{h}_{j}),$
where $\mathbf{h}_{i}$ and $\mathbf{h}_{j}$ are the updated features of case node $i$ and $j$ in the case pool $\mathcal{D}$. The matrix $\hat{\textbf{A}}\in\mathbb{R}^{n\times n}$ indicates a fully connected situation. And the degree regularisation is conducted on this pseudo adjacency matrix $\hat{\textbf{A}}$ only for candidate cases:
(11) $\ell_{\text{DegReg}}=\sum\limits^{o}_{i=1}\sum\limits^{n}_{j=1}(\hat{\textbf{A}}_{ij}),$